# Condor: A Code Discriminator Integrating General Semantics with Code Details

### v1

@@ 4.1.2. Effectiveness on Discrimination Scenario
Deepseek-Coder
Instruct (6.7B)
Meta-Llama-3.1
Instruct (8B)
Meta-Llama-3.1
Instruct (70B)
Original 43.23 36.81 52.64
Random 43.23 36.75 53.39
CodeRanker (110M) 42.84 37.48 57.69
CodeT5 (110M) 42.19 37.65 56.93
DeepSeek-Coder (1.3B) 43.54 37.32 58.90
Condor (110M) 43.73 39.60 59.64
Condor (1.3B) 50.85 45.27 62.63
Table 3 presents the performance of various discriminators on the CodeNanoFix dataset, evaluating the functional correctness of the top-1 code snippet using the Pass@1 metric after selecting outputs from the LLMs. The performance of the original generation models varies significantly, with Deepseek-Coder Instruct 6.7B achieving a Pass@1 of 43.23%, Meta-Llama-3.1 Instruct 8B of 36.81%, and Meta-Llama-3.1 Instruct 70B performing the best of 52.64%.
After applying a discriminator to the generated outputs, we observe that Condor (1.3B) significantly outperforms other discriminators. For example, Condor (1.3B) raises the Pass@1 score on Deepseek-Coder Instruct (6.7B) to 50.85%, resulting in a 17.63% increase compared to the original result. On Meta-Llama-3.1-Instruct (70B), Pass@1 reaches 62.63%, reflecting an improvement of nearly 19% over the original result. While Condor (110M) also shows improvements across multiple models, the gains are less pronounced compared to their larger counterparts, highlighting the importance of parameter size in capturing complex code details. In contrast, other discriminators (i.e., CodeRanker and CodeT5) also enhance the quality of generated results but yield smaller improvements, particularly in more complex generation tasks. The results of random selection show little difference from the original models, indicating that random selection has minimal impact on model performance.
{mdframed}
[linecolor=gray,roundcorner=12pt,backgroundcolor=gray!15,linewidth=3pt,innerleftmargin=2pt, leftmargin=0cm,rightmargin=0cm,topline=false,bottomline=false,rightline = false] Answer to RQ1: The experimental results demonstrate that code details play a crucial role in improving both classification and discrimination performance. By introducing code detail-aware strategies, Condor significantly enhances its ability to capture subtle code differences, leading to improved precision and recall across different models. For example, Condor (1.3B) achieves a precision of 74.39% and a recall of 72.40%. In the discrimination scenario, Condor’s performance is particularly outstanding. On Meta-Llama-3.1 Instruct 70B, Condor (1.3B) increases Pass@1 from 52.64% to 62.63%. This demonstrates Condor’s ability to accurately identify correct code and exhibit strong discriminative capabilities across models of various sizes.
4.2. RQ2: Generalizability to Code Generation Tasks
Tables 4 and 5 present the discrimination capabilities of our Condor model on the APPS and MBPP datasets, using the Pass@1 metric to measure Condor’s effectiveness in selecting LLM outputs. From these experimental results, we can observe Condor’s generalization ability in code generation tasks.
Table 4. Condor’s generalization capability in APPS.
Model’s Pass@1
Deepseek-Coder
Instruct (6.7B)
Meta-Llama-3.1
Instruct (8B)
Meta-Llama-3.1
Instruct (70B)
@@ 4.2. RQ2: Generalizability to Code Generation Tasks
CodeT5 (110M) 12.18 9.36 21.62
DeepSeek-Coder (1.3B) 12.10 9.64 22.42
Condor (110M) 12.24 10.32 22.32
@@
Table 5. Condor’s generalization capability in MBPP.
@@
Deepseek-Coder
Instruct (6.7B)
Meta-Llama-3.1
Instruct (8B)
Meta-Llama-3.1
Instruct (70B)
@@
CodeRanker (110M) 63.20 56.60 72.20
CodeT5 (110M) 61.60 58.20 72.80
DeepSeek-Coder (1.3B) 64.20 57.60 73.60
Condor (110M) 63.20 60.40 73.00
@@
In the APPS dataset (Table 4), due to the challenging nature of the algorithmic problems, the original generation models show relatively low performance. Deepseek-Coder Instruct 6.7B achieves a Pass@1 of 9.40%, Meta-Llama-3.1 Instruct 8B reaches 6.06%, and the Meta-Llama-3.1 Instruct 70B version achieves 10.16%. The results of random selection differ little from the original outputs, for example, on Deepseek-Coder Instruct 6.7B, Pass@1 only increases from 9.40% to 9.60%. In contrast, traditional discriminators like CodeRanker (110M) and CodeT5 (110M) show more significant improvements, with CodeRanker (110M) boosting Meta-Llama-3.1 Instruct 70B’s Pass@1 to 21.30%, and CodeT5 (110M) slightly higher at 21.62%. As the model size increases, DeepSeek-Coder (1.3B) further improves Pass@1 to 22.42%, demonstrating solid discriminative ability. Condor, enhanced by its code detail-awareness strategies, performs exceptionally well on the APPS dataset. Condor (110M) surpasses CodeRanker and CodeT5, increasing Meta-Llama-3.1 Instruct 8B’s Pass@1 to 10.32%, showcasing its discrimination advantage even in smaller models. Condor (1.3B) performs the best, achieving the highest Pass@1 across all models. For example, on the Meta-Llama-3.1-Instruct (70B), it boosts Pass@1 to 25.10%, representing an improvement over 147% from the original 10.16%. This demonstrates Condor’s strong ability to capture code details and significantly enhance performance in discrimination tasks.
@@
The experiments on the MBPP dataset (Table 5) further validate Condor’s generalization ability. Due to the relatively simpler nature of the MBPP problems, the original generation models’ Pass@1 scores are higher across the board. For example, Deepseek-Coder Instruct 6.7B achieves a Pass@1 of 59.00%, Meta-Llama-3.1 Instruct 8B reaches 54.00%, and Meta-Llama-3.1 Instruct 70B reaches 71.40%. Similar to the APPS dataset, random selection yields almost no improvement. CodeRanker (110M) and CodeT5 (110M) show more notable gains in discrimination performance, with CodeT5 (110M) improving Meta-Llama-3.1 Instruct 70B’s Pass@1 to 72.80%, slightly higher than CodeRanker (110M) at 72.20%. DeepSeek-Coder (1.3B) performs steadily on this dataset, raising Pass@1 to 73.60%. Similarly, Condor performs outstandingly on the MBPP dataset, especially Condor (1.3B), which further improves Meta-Llama-3.1 Instruct 70B’s Pass@1 to 75.20%, the best result. Condor (110M) also performs well on smaller models, boosting Meta-Llama-3.1 Instruct 8B’s Pass@1 to 60.40%, surpassing DeepSeek-Coder and other discriminators. Overall, Condor demonstrates competitiveness not only in smaller models but also in larger models, well utilizing its ability to capture code details and significantly improving Pass@1 accuracy.
{mdframed}
[linecolor=gray,roundcorner=12pt,backgroundcolor=gray!15,linewidth=3pt,innerleftmargin=2pt, leftmargin=0cm,rightmargin=0cm,topline=false,bottomline=false,rightline = false] Answer to RQ2: The experimental results demonstrate that Condor exhibits outstanding generalization capabilities, effectively improving discrimination accuracy on both the challenging APPS dataset and the relatively simpler MBPP dataset. On the APPS dataset, Condor (1.3B) increased Meta-Llama-3.1 Instruct 70B’s Pass@1 to 25.10%, compared to only 10.16% from the original generation. Similarly, on the MBPP dataset, Condor also performs exceptionally well, with Condor (1.3B) raising Meta-Llama-3.1 Instruct 70B’s Pass@1 to 75.20%, surpassing the original 71.40%. This indicates that Condor consistently enhances Pass@1 performance across tasks of various difficulty, showcasing its strong ability to capture subtle code details.
4.3. RQ3: Ablation Study
Table 6. Condor’s Ablation Study. The ‘w/o’ indicates without a specific strategy.
Model Precision Recall F1-Score
Pass@1 of DeepSeek-
Coder-Instruct (6.7B)
Pass@1 of Meta-Llama
-3.1-Instruct (8B)
Pass@1 of Meta-Llama
-3.1-Instruct (70B)
Condor (1.3B) 74.39 72.40 73.38 50.85 45.27 62.63
w/o Contrastive Learning 73.12 72.06 72.59 48.03 42.37 60.95
w/o Intermediate Data 67.98 74.66 71.16 45.29 38.54 59.84
@@ 4.3. RQ3: Ablation Study
Table 6 presents the ablation study results for our Condor model, analyzing performance changes on the CodeNanoFix dataset between the full model and its variants with specific strategies removed. Overall, the full Condor (1.3B) model delivers the best results across all metrics, with a Precision of 74.39%, Recall of 72.40%, and F1-score of 73.38%. It also excels in discrimination tasks, achieving a Pass@1 of 62.63% on Meta-Llama-3.1-Instruct (70B), highlighting its outstanding ability to capture code details and improve Pass@1 accuracy.
@@
When the embedding-level contrastive learning strategy is removed, precision drops to 73.12%, and recall also decreases to 72.06%, resulting in a lower F1-score of 72.59%. This indicates that contrastive learning plays a crucial role in enhancing the model’s ability to capture functional code differences. The discrimination performance also declines, with Pass@1 on Meta-Llama-3.1-Instruct (70B) dropping from 62.63% to 60.95%, demonstrating the significant contribution of contrastive learning to discrimination tasks. When the data-level intermediate data augmentation strategy is removed, the model shows more pronounced changes. Although recall increases slightly to 74.66%, suggesting the model becomes more conservative in recognizing correct code, precision drops significantly to 67.98%, leading to an overall F1-score decrease to 71.16%. The discriminative performance also declines, with Pass@1 on Meta-Llama-3.1-Instruct (70B) falling to 59.84%. This demonstrates the critical role of the intermediate data strategy in improving the model’s overall performance.
@@
Figure 3. Illustration of the impact of contrastive learning on code representations. The first three subplots show the 2D embeddings of code after training for 1 epoch (top left), 5 epochs (top right), and 40 epochs (bottom left). The bottom right subplot illustrates the changes in the average distance between correct code snippets and between error and correct code snippets as the number of training epochs increases.
@@
[linecolor=gray,roundcorner=12pt,backgroundcolor=gray!15,linewidth=3pt,innerleftmargin=2pt, leftmargin=0cm,rightmargin=0cm,topline=false,bottomline=false,rightline = false] Answer to RQ3: The ablation study results show that both the contrastive learning and intermediate data augmentation strategies effectively help the model capture detailed information in the code, significantly enhancing overall model performance. When contrastive learning is removed, both the model’s F1-score and its discrimination performance decline, indicating that contrastive learning is crucial for capturing functional code differences. Similarly, removing the intermediate data augmentation strategy slightly improves recall but significantly reduces precision by 8.6% and overall discrimination performance, highlighting this strategy’s crucial role in enhancing the model’s accuracy and discriminative ability.
5. Threats to Validity
@@
Threats to internal validity might come from the implementation of our experiments. To mitigate this threat, we use open-source models available on HuggingFace (Huggingface, 2024b) to minimize potential implementation errors. Additionally, to ensure the reproducibility of our experiments, we fix the random seed, ensuring consistency across different experimental results and reducing interference from randomness.
Threats to external validity may arise from two factors. The first factor may come from the choice of the base models. To mitigate this risk, we select the widely used and representative models CodeT5 and DeepSeek-Coder. CodeT5 is recognized as a classic encoder-decoder framework that has been validated and applied across multiple tasks, demonstrating strong versatility. DeepSeek-Coder is a decoder model that performs exceptionally well in a variety of software engineering tasks, providing a solid and reliable foundation for our experiments with our Condor model. The second potential threat comes from the downstream datasets used in our experiments. To mitigate this risk, we select the widely used MBPP and APPS datasets. These datasets are extensively used to evaluate code generation tasks and are commonly used to assess the performance of LLMs, providing strong support for the robustness of our experimental results.
@@
6. Related Work
6.1. Code Discriminators
@@
Existing approaches that improve code generation performance by selecting the outputs of LLMs can be categorized into two types: execution-based approaches and non-execution-based approaches. Execution-based approaches require executing the code and then using the feedback from the execution as additional information to enhance the discriminative effectiveness. Among them, MBR-EXEC, CodeT (Chen et al., 2023; Shi et al., 2022) requires executing the code to select the LLM’s outputs. However, in real-world development scenarios, it is not always easy to obtain the necessary test cases or testing environments for execution. Moreover, executing code generated by LLMs often raises safety concerns. Consequently, execution-based approaches face significant challenges in terms of flexibility. For flexibility, we focus on the non-execution-based approaches. Coder-Reviewer reranking (Zhang et al., 2023) uses prompt engineering by feeding the generated code to the original model to generate descriptions. It then selects the code by calculating the similarity between the generated description and the original description. However, as the current code model is quite large, making it costly to use the original model to assist in selecting generated outputs, which falls short of the efficiency requirements for real-world code generation tasks. Additionally, this approach is highly sensitive to prompt design and the model’s ability to interpret natural language, especially when problem descriptions are complex (e.g., APPS dataset). This discrepancy in the complexity of generating problem descriptions versus code can significantly impact the accuracy and consistency of the generated results. CodeRanker (Inala et al., 2022), which is similar to our work, improves model accuracy by incorporating a smaller classification model. Specifically, CodeRanker fine-tunes the CodeBERT model for code classification tasks, which in turn enhances the discrimination of the model’s outputs. Unlike existing models that simply leverage the inherent capabilities of models to represent code, Condor employs contrastive learning and data augmentation techniques to better detect subtle code details, thereby enhancing the discrimination performance.
6.2. Datasets for Code Discrimination

### v2

@@ IV-B RQ2: Generalizability to Code Generation Tasks
DS-Coder-Instruct
(6.7B, MP: 24.58)
Llama-3.1-Instruct
(8B, MP: 19.06)
Llama-3.1-Instruct
(70B, MP: 35.64)
@@
CodeT5 (110M) 12.18α 9.36 21.62
DS-Coder (1.3B) 12.10 9.64 22.42
DS-Coder-Instruct (1.3B, FS) 9.42 6.00 10.12
Llama-3.1-Instruct (8B, FS) 10.18 6.40 9.80
Condor (110M) 12.24α 10.32 22.32
@@
TABLE IV: Condor’s generalization capability in MBPP, where ‘FS’ indicates a few-shot manner and ‘MP’ represents the max potential.
@@
DS-Coder-Instruct
(6.7B, MP: 79.2)
Llama-3.1-Instruct
(8B, MP: 77.2)
Llama-3.1-Instruct
(70B, MP: 86.2)
@@
CodeRanker (110M) 63.20α 56.60 72.20
CodeT5 (110M) 61.60 58.20 72.80β
DS-Coder (1.3B) 64.20 57.60 73.60
DS-Coder-Instruct (1.3B, FS) 59.20 54.20 71.40
Llama-3.1-Instruct (8B, FS) 59.60 55.00 71.20
Condor (110M) 63.20α 60.40 73.00β
@@
TABLE V: Condor’s generalization capability in LiveCodeBench, where ‘FS’ indicates a few-shot manner and ‘MP’ represents the max potential.
Model’s Pass@1
DS-Coder-Instruct
(6.7B, MP: 18.86)
Llama-3.1-Instruct
(8B, MP: 15.11)
Llama-3.1-Instruct
(70B, MP: 35.11)
Original 15.68 12.73 28.97
Random 15.22 12.15 28.18
CodeRanker (110M) 15.11 12.50 29.09
CodeT5 (110M) 15.34α 12.50 29.43
DS-Coder (1.3B) 15.40 12.84β 29.68
DS-Coder-Instruct (1.3B, FS) 15.80 12.72 28.86
Llama-3.1-Instruct (8B, FS) 15.68 12.05 29.55
Condor (110M) 15.68α 13.07β 30.45
Condor (1.3B) 16.25 13.18β 30.90
@@
In the APPS dataset (Table III), due to the challenging nature of the algorithmic problems, the original generation models show relatively low performance. DS-Coder Instruct 6.7B achieves a Pass@1 of 9.40%, Llama-3.1 Instruct 8B reaches 6.06%, and the Llama-3.1 Instruct 70B version achieves 10.16%. The results of random selection differ little from the original outputs, for example, on DS-Coder-Instruct 6.7B, Pass@1 only increases from 9.40% to 9.60%. In contrast, traditional discriminators like CodeRanker (110M) and CodeT5 (110M) show more significant improvements, with CodeRanker (110M) boosting Llama-3.1-Instruct 70B’s Pass@1 to 21.30%, and CodeT5 (110M) slightly higher at 21.62%. As the model size increases, DS-Coder (1.3B) further improves Pass@1 to 22.42%, demonstrating solid discriminative ability. Condor, enhanced by its code detail-awareness strategies, performs exceptionally well on the APPS dataset. Condor (110M) surpasses CodeRanker and CodeT5, increasing Llama-3.1 Instruct 8B’s Pass@1 to 10.32%, showcasing its discrimination advantage even in smaller models. Condor (1.3B) performs the best, achieving the highest Pass@1 across all models. For example, on the Llama-3.1-Instruct (70B), it boosts Pass@1 to 25.10%, representing an improvement over 147% from the original 10.16%. This demonstrates Condor’s strong ability to capture code details and significantly enhance performance in discrimination tasks.
Figure 3: Case studies showing Condor’s ability to recognize correct solutions and detect errors in code.
@@
The experiments on the MBPP dataset (Table IV) further validate Condor’s generalization ability. Due to the relatively simpler nature of the MBPP problems, the original generation models’ Pass@1 scores are higher across the board. For example, DS-Coder Instruct 6.7B achieves a Pass@1 of 59.00%, Llama-3.1-Instruct 8B reaches 54.00%, and Llama-3.1-Instruct 70B reaches 71.40%. Similar to the APPS dataset, random selection yields almost no improvement. CodeRanker (110M) and CodeT5 (110M) show more notable gains in discrimination performance, with CodeT5 (110M) improving Llama-3.1 Instruct 70B’s Pass@1 to 72.80%, slightly higher than CodeRanker (110M) at 72.20%. DS-Coder (1.3B) performs steadily on this dataset, raising Pass@1 to 73.60%. Similarly, Condor performs outstandingly on the MBPP dataset, especially Condor (1.3B), which further improves Llama-3.1 Instruct 70B’s Pass@1 to 75.20%, the best result. Condor (110M) also performs well on smaller models, boosting Llama-3.1 Instruct 8B’s Pass@1 to 60.40%, surpassing DS-Coder and other discriminators. Overall, Condor demonstrates competitiveness not only in smaller models but also in larger models, well utilizing its ability to capture code details and significantly improving Pass@1 accuracy.
Figure 3 presents two cases in MBPP dataset, highlighting Condor’s discrimination abilities at different granularity levels. In the first example (top, Task-86), Condor distinguishes between two formulas for the centered hexagonal number and correctly identifies the accurate computation based on fine-grained numerical differences. In the lower example (bottom, Task-344), the task requires counting the number of elements with odd factors in a given range. Condor is faced with a more challenging scenario involving logic-level reasoning. The first candidate solution checks for perfect squares and then further filters those whose square roots are odd, which does not align with the problem’s requirement of counting all perfect squares (since every perfect square has an odd number of factors, regardless of the parity of the square root). Condor successfully recognizes the discrepancy in the logical process and selects the correct solution, showing its capability to capture not only numerical but also coarse-grained semantic differences. These cases highlight that Condor can accurately discriminate between candidates based on both precise computational correctness and broader algorithmic logic, supporting its robust performance across a variety of code reasoning scenarios.
Following our analysis of the APPS and MBPP, we further evaluate Condor on LiveCodeBench (Table V), a challenging new benchmark that emphasizes dynamic and real-world code tasks. Importantly, LiveCodeBench does not provide a dedicated training set for discriminators, making it more challenging for evaluating true model generalization and real-world applicability. The original Pass@1 scores are moderate (e.g., DS-Coder-Instruct 6.7B: 15.68%, Llama-3.1-Instruct 8B: 12.73%, Llama-3.1-Instruct 70B: 28.97%), with random selection again yielding almost no gain. The improvements brought by traditional discriminators are modest, and their effects are even less pronounced for this more complex dataset. Instead, Condor continues to show a strong generalization advantage. For example, Condor (1.3B) pushes Llama-3.1-Instruct 70B’s Pass@1 to 30.90%, the highest among all tested discriminators, while smaller-scale Condor models also surpass the other baselines on various settings. Although the gains on LiveCodeBench are smaller than on APPS, this is because the maximum potential (MP) for improvement is lower, as shown in these tables (i.e., ‘MP’ in Table IIIIVV). With fewer correct candidates generated, even the best discriminator can provide only limited gains, highlighting the challenge of this benchmark.
Answer to RQ2: The experimental results demonstrate that Condor exhibits outstanding generalization capabilities, effectively improving discrimination accuracy on the APPS, MBPP, and LiveCodeBench datasets. On APPS, Condor (1.3B) raises Llama-3.1-Instruct 70B’s Pass@1 from 10.16% to 25.10%. On MBPP, Condor (1.3B) boosts Pass@1 from 71.40% to 75.20%. Similarly, on LiveCodeBench, Condor consistently delivers the highest Pass@1 among all tested discriminators, further validating its robustness and effectiveness across diverse and challenging benchmarks.
IV-C RQ3: Ablation Study
@@
The lower part of Table II presents the ablation study results for our Condor model, comparing the full model to its variants with individual strategies removed. We use the 1.3B version of Condor for ablation because it achieves the most balanced and robust performance across all metrics, with a precision of 74.39%, recall of 72.40%, and F1-score of 73.38%. In discrimination tasks, Condor (1.3B) also achieves a leading Pass@1 of 62.63% on Llama-3.1-Instruct (70B), underscoring its effectiveness in capturing code details and serving as a strong baseline for analyzing the impact of each strategy.
@@
When the embedding-level contrastive learning strategy is removed, precision drops to 73.12%, and recall also decreases to 72.06%, resulting in a lower F1-score of 72.59%. This indicates that contrastive learning plays a crucial role in enhancing the model’s ability to capture functional code differences. The discrimination performance also declines, with Pass@1 on Llama-3.1-Instruct (70B) dropping from 62.63% to 60.95%, demonstrating the significant contribution of contrastive learning to discrimination tasks. When the data-level intermediate data augmentation strategy is removed, the model shows more pronounced changes. Although recall increases slightly to 74.66%, suggesting the model becomes more conservative in recognizing correct code, precision drops significantly to 67.98%, leading to an overall F1-score decrease to 71.16%. The discriminative performance also declines, with Pass@1 on Llama-3.1-Instruct (70B) falling to 59.84%. This demonstrates the critical role of the intermediate data strategy in improving the model’s overall performance.
@@
Figure 4: Illustration of the impact of contrastive learning on code representations. The first three subplots show the 2D embeddings of code after training for 1 epoch (top left), 5 epochs (top right), and 40 epochs (bottom left). The bottom right subplot illustrates the changes in the average distance between correct code snippets and between error and correct code snippets as the number of training epochs increases.
@@
Answer to RQ3: The ablation study results show that both the contrastive learning and intermediate data augmentation strategies effectively help the model capture detailed information in the code, significantly enhancing overall model performance. When contrastive learning is removed, both the model’s F1-score and its discrimination performance decline, indicating that contrastive learning is crucial for capturing functional code differences. Similarly, removing the intermediate data augmentation strategy slightly improves recall but significantly reduces precision by 8.6% and overall discrimination performance, highlighting this strategy’s crucial role in enhancing the model’s accuracy and discriminative ability.
V Threats to Validity
@@
Threats to internal validity might come from the implementation of our experiments. To mitigate this threat, we use open-source models available on HuggingFace [huggingface] to minimize potential implementation errors. Additionally, to ensure the reproducibility of our experiments, we fix the random seed, ensuring consistency across different experimental results and reducing interference from randomness.
Threats to external validity may arise from fourth factors. The first factor may come from the choice of the base models. To mitigate this risk, we select the widely used and representative models CodeT5 and DS-Coder. CodeT5 is recognized as a classic encoder-decoder framework that has been validated and applied across multiple tasks, demonstrating strong versatility. DS-Coder is a decoder model that performs exceptionally well in a variety of software engineering tasks, providing a solid and reliable foundation for our experiments with our Condor model. The second potential threat may arise from the choice of downstream datasets in our experiments. To mitigate this risk, we select the widely used code generation datasets, such as APPS, MBPP, and LiveCodeBench. These datasets are extensively used to evaluate code generation tasks and are commonly used to assess the performance of LLMs, providing strong support for the robustness of our experimental results. In the future, we plan to extend our evaluation to datasets beyond code generation and code repair tasks, such as ChangeGuard [groninger2025changeguard], which focuses on functional change detection in code edits. The third potential threat may come from the choice of programming language. Our current experiments are conducted exclusively on Python, which, while widely used in AI research and competitive programming, does not fully represent the diversity of programming languages used in real-world software engineering (e.g., Java, C++, JavaScript, Go, Rust). To mitigate this, we design Condor’s architecture to be language-agnostic, allowing it to be adapted to other languages by collecting appropriate training data and re-tokenizing. Exploring multi-language settings remains an important direction for future work. The fourth potential threat might arise from the gap between benchmark settings and real-world development environments. In large-scale software engineering, correctness often depends on the broader system context, including dependencies and cross-module interactions, which are not captured when evaluating isolated code snippets. To mitigate this gap, we introduce LiveCodeBench into our evaluation. LiveCodeBench contains real-world, dynamic coding tasks without a public training set, making it a strict test of model generalization beyond the training distribution. Moreover, our results show that Condor scales well—larger versions (e.g., 1.3B parameters) achieve more promising performance gains, indicating potential for further improvements when moving towards more complex, production-scale scenarios.
@@
VI Related Work
VI-A Code Discriminators
@@
Existing approaches that improve code generation performance by selecting the outputs of LLMs can be categorized into two types: execution-based approaches and non-execution-based approaches. Execution-based approaches require executing the code and then using the feedback from the execution as additional information to enhance the discriminative effectiveness. Among them, MBR-EXEC, CodeT [codet, mbrexec] requires executing the code to select the LLM’s outputs. However, in real-world development scenarios, it is not always easy to obtain the necessary test cases or testing environments for execution. Moreover, executing code generated by LLMs often raises safety concerns. Consequently, execution-based approaches face significant challenges in terms of flexibility. For flexibility, we focus on the non-execution-based approaches. Coder-Reviewer reranking [coder_reviewer_reranking] employs prompt engineering by feeding the generated code back into the original model to generate descriptions, and then selects candidate code based on the similarity between the generated and original descriptions. However, as current code models are quite large, using the original model for reranking is computationally expensive and fails to meet the efficiency requirements of real-world code generation tasks. Furthermore, this approach is highly sensitive to prompt design and the model’s ability to interpret natural language, especially when problem descriptions are complex (e.g., APPS dataset). The gap between the complexity of generating problem descriptions and generating code can lead to significant variability in the accuracy and consistency of the reranking results. CodeRanker [coderanker], which is similar to our work, improves model accuracy by incorporating a smaller classification model. Specifically, CodeRanker fine-tunes the CodeBERT model for code classification tasks, which in turn enhances the discrimination of the model’s outputs. Unlike these approaches, Condor adopts a code-centric perspective. Instead of relying on natural language interpretation or simple code representations, Condor leverages contrastive learning and data augmentation to directly analyze the structural and semantic differences between code candidates. This enables Condor to better detect subtle variations in code details that may significantly impact program functionality, thereby providing more robust and efficient discrimination of correct solutions.
VI-B Datasets for Code Discrimination